# Double white dwarf binary population in MOCCA star clusters

### v2

@@
(Accepted XXX. Received YYY; in original form ZZZ)
There could be a significant population of double white dwarf binaries (DWDs) inside globular clusters (GCs), however, these are often too faint to be individually observed. We have utilized a large number GC models evolved with the Monte Carlo Cluster Simulator (MOCCA) code, to create a large statistical dataset of DWDs. These models include multiple-stellar populations, resulting in two distinct initial populations: one dense and another less dense. Due to the lower density of one population, a large number of objects escape during the early GC evolution, leading to a high mass-loss rate. In this dataset we have analysed three main groups of DWDs, namely in-cluster binaries, escaped binaries, and isolated evolution of primordial binaries. We compared the properties of these groups to observations of close and wide binaries. We find that the number of escaping DWDs is significantly larger than the number of in-cluster binaries and those that form via the isolated evolution of all promiridial binaries in our GC models. This suggests that dynamics play an important role in the formation of DWDs. For close binaries, we found a good agreement in the separations of escaped binaries and isolated binaries, but in-cluster binaries showed slight differences. We could not reproduce the observed extremely low mass WDs due to the limitations of our stellar and binary evolution prescriptions. For wide binaries, we also found a good agreement in the separations and masses, after accounting for observational selection effects. We conclude that, even though the current observational samples of DWDs are extremely biased and incomplete, our results compare reasonably well with observations.
@@ 1 Introduction
Globular clusters (GCs) are self-gravitating spherical collections of tens of thousands to up to several millions of stars. The cores of these clusters are very dense (up to ${\sim 10^{6}}$ stars per pc3), making stellar collisions and dynamical interactions relatively frequent. These interactions can affect not only the stars but also the cluster as a whole. For a binary, an interaction with another star can lead to a hardening or softening, depending on the interaction. It can also lead to the formation of a whole new binary through an exchange or a merger in a three-body or four-body interaction. In some cases, these interactions can even disrupt the binary so that it dissolves. During a three-body interaction, binaries can increase the kinetic energy of a third object or of the binary’s centre of mass.
@@
Globular clusters are generally very old and without any on-going star formation. This causes them to almost exclusively contain old stars and, combined with the fact that they are found in almost all galaxies, makes them very interesting targets for observations and theoretical studies. Because of this, several research groups around the world are working on codes to simulate the evolution of these clusters, in order to better understand their evolution. Some examples of these codes include but are not limited to the mocca code (Giersz et al. 2013), cmc (Kremer et al. 2020; Rodriguez et al. 2021), nbody6++GPU (Wang et al. 2015, 2016; Kamlah et al. 2022; Arca Sedda et al. 2024), nbody7 (Banerjee 2022) and PeTar (Wang et al. 2020).
@@
White dwarfs (WDs) are relatively abundant in GCs (Torres et al. 2015) and often make up a large portion of the stellar population. White dwarfs are important objects inside clusters since they can, among other things, be used to determine the age of their host cluster (e.g. Fontaine et al. 2001). The cooling process of WDs has been believed to be quite well understood so observations of their temperature and luminosity can be used to determine the age of the WD and thus the cluster. However, recent developments in understanding the physical processes occurring during their cooling have been challenging this picture (e.g. Bédard et al. 2024)
@@
There are several different populations of WDs. The first and most basic one is that composed of single WDs. Another sizeable population is made up of detached binaries hosting main-sequence (MS) stars. In this case, the WD and the MS star have a wide enough orbit so that each star does not fill its Roche lobe and thus there is no ongoing mass transfer Campos et al. (2018). The semi-detached counterpart of these systems is called cataclysmic variable. A cataclysmic variable is a binary with a WD and a Roche-lobe filling secondary star, typically a MS star. These binaries are very tight with a short period (${\sim 1-6}$ h) such that the secondary fills its Roche lobe and transfers mass onto the WD. This mass transfer typically results in an accretion disk around the WD which often can be seen in ultraviolet and X-ray observations. The last population of WDs we consider are the focus of this paper and corresponds to double WD binaries (DWDs) which consist of two WDs.
@@
Double WD binaries can, in general, form in two ways: the progenitors to the WDs were formed together in a binary system (Belloni & Schreiber 2023a) or the binary was formed during a dynamical interaction (Kratter 2011). The resulting DWD can have different properties depending on the formation mechanism. Properties of primordial DWDs, that is, DWDs forming from the evolution of primordial binaries, depend on the initial binary parameter distribution of primordial binaries. The eccentricity distribution of a population of primordial binaries will have a thermal distribution while a population of dynamically formed binaries will have higher eccentricities than a thermal distribution. Meanwhile, the semi-major axis distribution of primordial binaries will depend on the initial distributions while for dynamically formed binaries it depends on the properties of the dense environment and gravitational encounters in which the binary is formed. In addition to this, due to the dense environments where DWDs are formed, it is likely that the binary can eventually experience more dynamical interactions which may affect the semi-major axis further.
@@
Recent studies have indicated that massive WDs most likely dominate the inner-most regions of core-collapsed GCs such as NGC 6397 (Kremer et al. 2021). These populations are, through dynamical interactions, stopping the cluster from completely collapsing. This produces very tight, inspiraling WD binaries that may be observed with future telescopes.
This paper will focus on DWDs from simulations of GCs with multiple-stellar populations (Lee et al. 1999; Gratton et al. 2012; Bastian & Lardo 2018). We will focus on in-cluster, escaped and isolated binaries (i.e. binaries with the same properties as the ones inside the clusters but evolved in isolation without dynamical interactions or relaxation affecting them). These binaries are then compared to observations of DWDs in two regions; close binaries where the data is retrieved from Brown et al. (2020), Kosakowski et al. (2023) and Schreiber et al. (2022) and wide binaries where the data is retrieved from El-Badry et al. (2021) and Heintz et al. (2022). This was done to get a better understanding of how realistic the DWD populations are in clusters from our simulations. We also wanted to find out if there are differences between the DWD populations inside and outside of clusters (escapers and binaries evolved in isolation). Finding any potential differences would be an important discovery since it could aid with determining the origin of observed WD binaries. This also relates to an upcoming study where we will investigate the GW signals from WD binaries, if we would find differences in these populations then we can also expect to see differences in their respective GW signals. This comparison, however, proved to be more difficult than we expected, due to severe limitations in the parameter space of observations, we have to limit our data greatly in order to make a proper comparison. Nevertheless, when restricting our data in ways that agree with observational techniques and limitations we find good agreements everywhere except for the masses of close DWDs.
@@ 2.1 Close binaries
The data for close DWDs are obtained from Brown et al. (2020), Kosakowski et al. (2023) and Schreiber et al. (2022). Brown et al. (2020) uses the Extremely low mass (ELM) survey to find a sample of 98 detached, close double WD binaries. This survey targeted $<0.3$ M helium-core WDs binaries with periods between $0.0089$ and 1.5 day. The absolute magnitude in the G band of the faintest object in this sample is $11.24$ mag. These WDs are so close to each other that they cannot be resolved separately. Therefore this magnitude limit is for both WDs combined. Kosakowski et al. (2023) is a continuation of the ELM survey where they found 28 additional DWDs. Schreiber et al. (2022) collected 57 close DWDs from different surveys and did not limit themselves to ELMs.
@@ 2.2 Wide binaries
For wide binaries we use the published catalogue from Heintz et al. (2022) which contains data of 1590 wide DWDs from Gaia EDR3 (Gaia Collaboration et al. 2021) and El-Badry et al. (2021). The goal of Heintz et al. (2022) was to check existing methods for determining WD ages. They used wide DWDs, separated by more than 100 au, and made the assumption that at these distances, the stars would evolve independently without any significant interactions between the two binary components. In addition, they assumed that the stars in the binary were born from the same molecular cloud at approximately the same time.
This data set includes a chance alignment factor ($R\_chance\_align$) which determines the chance that the detected binary is in fact not a binary but two single stars that, due to chance, appear to be bound in a binary. This produces an upturn in the tail at larger separations. Section 3 of El-Badry et al. (2021) explains this in detail. We use this chance alignment factor and remove all binaries with $R\_chance\_align>0.1$ which leaves us with 1389 binaries.
The apparent magnitudes of the WDs in the Gaia G-band, $G$, are limited to 13 mag ¡ G ¡ 21 mag, however, this is not a cut done by the authors but a limitation on observational instruments. Using the parallax, $p$, we can estimate the absolute magnitude with
$G_{abs}=G+5-5\log\left(\frac{1}{p}\right)$ (1)
Using the parallaxes and apparent G-band magnitudes from Heintz et al. (2022) we estimate that the faintest WD in this data set will have an absolute magnitude in the G band of $\sim$16 mag. Thus we will remove all binaries where any member has a magnitude above this limit from our numerical data.
Figure 7 in Heintz et al. (2022) shows that in their dataset, for a large number of binaries, the cooling age of the primary, more massive, star is shorter than the secondary. This is not expected in isolated binaries and might indicate that these systems used to be triple systems where the inner binary interacted through mass transfer or a merger and later this inner binary was split or merged and thus formed a binary with the third object. These objects are removed from our comparisons.
@@ 3.1 The mocca code
We use the MOnte Carlo Cluster simulAtor, mocca  [^1] http://moccacode.net, (Hypki & Giersz 2013; Giersz et al. 2013) and more specifically the simulations from MOCCA-SURVEY Database III which is an upgrade over the MOCCA-SURVEY Database II (Hypki et al. 2022) and MOCCA-SURVEY Database I (Askar et al. 2017). mocca is a GC simulator built on a Monte Carlo approach. It allows the usage of multiple stellar populations with different properties, large control over initial cluster conditions and allows quick evolution of large clusters with more than 1 million initial members.
@@
Since the publication of bse, there have been several upgrades to the code (see, e.g. Kamlah et al. 2022, for more details). The main upgrades that are relevant for this study are the inclusion of a proper prescription for cataclysmic variable evolution (Belloni et al. 2018b, 2019) and the inclusion of improved wind prescriptions (Belczynski et al. 2016).
There are though two main problems in bse which play an important role on the types of binaries we are interested in. The first one is related to WD evolution, which is not properly handled by bse. The luminosity evolution of WDs is modelled in bse using the Mestel (1952) standard cooling theory. This theory though does not take into account important properties such as core composition and surface abundances as well as crucial features such as crystallization, which can significantly delay the WD cooling for Gyr and has been shown to be very important for explaining several features in WD populations (e.g. Tremblay et al. 2019; Schreiber et al. 2021; Bagnulo & Landstreet 2022; Blatman & Ginzburg 2024)
We therefore improved WD evolution in our analysis by computing more accurate WD properties via interpolation/extrapolation using the cooling sequences calculated by Bédard et al. (2020), assuming a thin hydrogen atmosphere. Provided the WD age and mass, we linearly interpolate/extrapolate through the sequences to obtain the WD properties such as effective temperature, radius and luminosity. These sequences were computed assuming solar metallicity, which is higher than assumed in our GC simulations. However, the impact of metallicity on WD evolution regarding the properties in which we are interested is virtually negligible (Renedo et al. 2010). That said, for the purposes of our analysis, it is reasonable to use the solar-metallicity sequences calculated by Bédard et al. (2020). The main motivation to use such sequences amongst the few that are available comes from the fact that they were used by Heintz et al. (2022).
@@ 3.2 Initial conditions
We use 197 cluster simulations with a wide range of initial cluster parameters. A full description of the different cluster initial conditions can be found in Hypki et al. (2024, in prep). The most important parameters are:
  • We have many different combinations of sizes of clusters, from 550 thousand to 2.6 million initial objects (two multiple stellar populations).
  • We have both tidally filling and underfilling clusters
  • We have both 10% and 95% initial binary fraction
  • We use a Kroupa mass function between 0.08 and 150 M
  • Different Galactocentric distance
  • Different properties of multiple populations
A big difference with these simulations compared to earlier simulations is that we are using multiple-stellar populations in the framework of asymptotic giant branch scenario (Calura et al. 2019) in these new simulations and thus a new path to GC evolution. Instead of having one tidally underfilling dense cluster we have two populations; one tidally filling (or slightly tidally underfilling), less dense population with a higher maximum initial stellar mass and one tidally underfilling, denser population. This leads to a large number of early escapers from the first population where approximately 30-40% of mass is removed in the first few Myr. These binaries are mostly primordial and undisturbed from dynamical interactions since they escape at an early stage before they have had time to interact with other objects. The second population is more dense but the maximum stellar mass is lower, there is no formation of massive stars in this population. This limit is set to 20 M which means that there are no supernovae forming BHs from the second population. The previously used initial cluster conditions are quite similar to the second population while the first population is not as tightly bound and is able to lose a lot of mass early in the evolution.
The choice of 10% or 95% binary fraction affects the initial binary properties. For example, in a 10% binary fraction cluster, the largest semi-major axis value is 100 au and the distribution is uniform in log(a). With 95% binary fraction, we allow binaries to form with larger separations and the initial distribution of semi-major axis and eccentricity is based on Kroupa (1995); Belloni et al. (2017). Since these binaries are often disrupted by dynamical interactions, a 95% binary fraction has been found to agree well with observations (Belloni et al. 2017, 2018a). For both binary fractions, the initial eccentricity distribution is thermal.

### v3

@@
(Received 8 May 2024 / Accepted 18 July 2024)
There could be a significant population of double white dwarf binaries (DWDs) inside globular clusters (GCs); however, these binaries are often too faint to be individually observed. We have utilized a large number GC models evolved with the Monte Carlo Cluster Simulator (MOCCA) code to create a large statistical dataset of DWDs. These models include multiple-stellar populations, resulting in two distinct initial populations: one dense and the other less dense. Due to the lower density of one population, a large number of objects escape during the early GC evolution, leading to a high mass-loss rate. In this dataset we have analyzed three main groups of DWDs, namely in-cluster binaries, escaped binaries, and binaries formed from the isolated evolution of primordial binaries. We compared the properties of these groups to observations of close and wide binaries. We find that the number of escaping DWDs is significantly larger than the number of in-cluster binaries and those that form via the isolated evolution of all primordial binaries in our GC models. This suggests that dynamics play an important role in the formation of DWDs. For close binaries, we found a good agreement in the separations of escaped binaries and isolated binaries, but in-cluster binaries showed slight differences. We could not reproduce the observed extremely low mass WDs due to the limitations of our stellar and binary evolution prescriptions. For wide binaries, we also found a good agreement in the separations and masses, after accounting for observational selection effects. Even though the current observational samples of DWDs are extremely biased and incomplete, we conclude that our results compare reasonably well with observations.
@@ 1 Introduction
Globular clusters (GCs) are self-gravitating spherical collections of tens of thousands of stars up to several million stars. The cores of these clusters are very dense (up to ${\sim 10^{6}}$ stars per pc3), making stellar collisions and dynamical interactions relatively frequent. These interactions can affect not only the stars, but also the cluster as a whole. For a binary, an interaction with another star can lead to a hardening or softening, depending on the interaction. It can also lead to the formation of a whole new binary through an exchange or a merger in a three-body or four-body interaction. In some cases these interactions can even disrupt the binary so that it dissolves. During a three-body interaction, binaries can increase the kinetic energy of a third object or of the binary’s center of mass.
@@
Globular clusters are generally very old and without any ongoing star formation. This causes them to almost exclusively contain old stars and, combined with the fact that they are found in almost all galaxies, makes them very interesting targets for observations and theoretical studies. Because of this, several research groups around the world are working on codes to simulate the evolution of these clusters in order to better understand their evolution. Some examples of these codes include, but are not limited to, the MOCCA code (Giersz et al. 2013), CMC (Kremer et al. 2020; Rodriguez et al. 2021), nbody6++GPU (Wang et al. 2015, 2016; Kamlah et al. 2022; Arca Sedda et al. 2024), nbody7 (Banerjee 2022), and PeTar (Wang et al. 2020).
@@
White dwarfs (WDs) are relatively abundant in GCs (Torres et al. 2015), and often make up a large portion of the stellar population. White dwarfs are important objects inside clusters since they can, among other things, be used to determine the age of their host cluster (e.g., Fontaine et al. 2001). The cooling process of WDs is believed to be quite well understood, so observations of their temperature and luminosity can be used to determine the age of the WD, and thus the cluster. However, recent developments in understanding the physical processes occurring during their cooling have been challenging this picture (e.g., Bédard et al. 2024)
@@
There are several different populations of WDs. The first and most basic one is that composed of single WDs. Another sizable population is made up of detached binaries hosting main-sequence (MS) stars. In this case, the WD and the MS star have a wide enough orbit so that each star does not fill its Roche lobe, and thus there is no ongoing mass transfer Campos et al. (2018). The semi-detached counterpart of these systems is called a cataclysmic variable. A cataclysmic variable is a binary with a WD and a Roche-lobe filling secondary star, typically a MS star. These binaries are very tight with a short period (${\sim 1-6}$ h) such that the secondary fills its Roche lobe and transfers mass onto the WD. This mass transfer typically results in an accretion disk around the WD, which can often be seen in ultraviolet and X-ray observations. The last population of WDs we consider are the focus of this paper, and correspond to double WD binaries (DWDs), which consist of two WDs.
@@
Double WD binaries can, in general, form in two ways: the progenitors to the WDs were formed together in a binary system (Belloni & Schreiber 2023a) or the binary was formed during a dynamical interaction (Kratter 2011). The resulting DWD can have different properties depending on the formation mechanism. The properties of primordial DWDs, that is DWDs forming from the evolution of primordial binaries, depend on the initial binary parameter distribution of primordial binaries. The eccentricity distribution of a population of primordial binaries, by definition and according to our setup, has a thermal distribution (Jeans 1919), while a population of dynamically formed binaries has higher eccentricities than a thermal distribution since they experience only a few dynamical interactions that will, on average, increase their eccentricities. Meanwhile, the semi-major axis distribution of primordial binaries depends on the initial distributions, while for dynamically formed binaries it depends on the properties of the dense environment and gravitational encounters in which the binary is formed. In addition, due to the dense environments where DWDs are formed, it is likely that the binary can eventually experience more dynamical interactions that may affect the semimajor axis further.
@@
Recent studies have indicated that massive WDs most likely dominate the innermost regions of core-collapsed GCs such as NGC 6397 (Kremer et al. 2021). These populations, through dynamical interactions, stop the cluster from completely collapsing. This produces very tight, inspiraling WD binaries that may be observed with future telescopes.
This paper focuses on DWDs from simulations of GCs with multiple-stellar populations (Lee et al. 1999; Gratton et al. 2012; Bastian & Lardo 2018). We focus on in-cluster, escaped, and isolated binaries (i.e., binaries with the same properties as those inside the clusters, but evolved in isolation without dynamical interactions or relaxation affecting them). These binaries are then compared to observations of DWDs in two regions; close binaries where the data is retrieved from Brown et al. (2020), Kosakowski et al. (2023), and Schreiber et al. (2022) and wide binaries where the data is retrieved from El-Badry et al. (2021) and Heintz et al. (2022). This was done to get a better understanding of how realistic the DWD populations are in clusters from our simulations. We also wanted to find out if there are differences between the DWD populations inside and outside of clusters (escapers and binaries evolved in isolation). Finding any potential differences would be an important discovery since it could aid in determining the origin of observed WD binaries and help our understanding of whether the field population of DWDs can be attributed to escapers from clusters. This also relates to an upcoming study where we will investigate the GW signals from WD binaries: if we find differences in these populations, then we can also expect to see differences in their respective GW signals. This comparison, however, proved to be more difficult than we expected; due to severe limitations in the parameter space of observations, we had to limit our data greatly in order to make a proper comparison. Nevertheless, when restricting our data in ways that agree with observational techniques and limitations we find good agreement everywhere, except for the masses of close DWDs.
@@ 2.1 Close binaries
The data for close DWDs were obtained from Brown et al. (2020), Kosakowski et al. (2023), and Schreiber et al. (2022). Brown et al. (2020) uses the Extremely Low Mass (ELM) Survey to find a sample of 98 detached, close double WD binaries. This survey targeted $<0.3$ M helium-core WDs binaries with periods between $0.0089$ and 1.5 days. The absolute magnitude in the G band of the faintest object in this sample is $11.24$ mag. These WDs are so close to each other that they cannot be resolved separately. Therefore, this magnitude limit is for the two WDs combined. Kosakowski et al. (2023) is a continuation of the ELM Survey where they found 28 additional DWDs. One important point to mention is that ELM WDs are rare; O’Brien et al. (2024) found 54 candidate DWDs, but none of them can be considered ELMs. Unfortunately, O’Brien et al. (2024) does not include periods for any of these close binaries. Instead, we add another sample of higher mass, close DWDs from Schreiber et al. (2022) where 57 close DWDs were collected from different surveys and there was no limitation to only ELMs.
@@ 2.2 Wide binaries
For wide binaries we used the published catalog from Heintz et al. (2022), which contains data of 1590 wide DWDs from Gaia EDR3 (Gaia Collaboration et al. 2021) and El-Badry et al. (2021). The goal of Heintz et al. (2022) was to check existing methods for determining WD ages. They used wide DWDs, separated by more than 100 au, and made the assumption that at these distances, the stars would evolve independently without any significant interactions between the two binary components. In addition, they assumed that the stars in the binary were born from the same molecular cloud at approximately the same time.
This dataset includes a chance alignment factor ($R\_chance\_align$), which determines the chance that the detected binary is in fact not a binary but two single stars that, due to chance, appear to be bound in a binary. This produces an upturn in the tail at larger separations. Section 3 of El-Badry et al. (2021) explains this in detail. We use this chance alignment factor and remove all binaries with $R\_chance\_align>0.1$, which leaves us with 1389 binaries.
The apparent magnitudes of the WDs in the Gaia G band, $G$, are limited to 13 mag ¡ G ¡ 21 mag; however, this is not a cut done by the authors, but a limitation on observational instruments. Using the parallax, $p$, we can estimate the absolute magnitude with
$G_{abs}=G+5-5\log\left(\frac{1}{p}\right).$ (1)
Using the parallaxes and apparent G-band magnitudes from Heintz et al. (2022) we estimate that the faintest WD in this dataset will have an absolute magnitude in the G band of $\sim$16 mag. In addition to this, due to the finite age of the Galactic disk, there is a peak in the white dwarf luminosity function at $\sim$16 mag (García-Berro & Oswalt 2016). Thus, we remove from our numerical data all binaries where either member has a magnitude above this limit.
Figure 7 in Heintz et al. (2022) shows that in their dataset, for a large number of binaries, the cooling age of the primary, the more massive star, is shorter than the secondary. This is not expected in isolated binaries, and might indicate that these systems used to be triple systems where the inner binary interacted through mass transfer or a merger, and that this inner binary was later split or merged, and thus formed a binary with the third object. It is also possible that a nonmonotonic initial-to-final mass relation can cause this (Marigo et al. 2020). These objects were removed from our comparisons.
@@ 3.1 The MOCCA code
We use the MOnte Carlo Cluster simulAtor, MOCCA  [^1] http://moccacode.net (Hypki & Giersz 2013; Giersz et al. 2013), and more specifically the simulations from MOCCA-SURVEY Database III (Hypki et al. 2024), which is an upgrade over the MOCCA-SURVEY Database II (Hypki et al. 2022) and MOCCA-SURVEY Database I (Askar et al. 2017). MOCCA is a GC simulator built on a Monte Carlo approach. It allows the usage of multiple stellar populations with different properties, provides greater control over initial cluster conditions, and allows the quick evolution of large clusters with more than one million initial members.
@@
Since the publication of BSE, there have been several upgrades to the code (see, e.g., Kamlah et al. 2022, for more details). The main upgrades that are relevant for this study are the inclusion of a proper prescription for cataclysmic variable evolution (Belloni et al. 2018b, 2019) and the inclusion of improved wind prescriptions (Belczynski et al. 2016).
There are, however, two main problems in BSE that play an important role in the types of binaries we are interested in. The first is related to WD evolution, which is not properly handled by BSE. The luminosity evolution of WDs is modeled in BSE using the Mestel (1952) standard cooling theory. This theory though does not take into account important properties such as core composition and surface abundances as well as crucial features such as crystallization, which can significantly delay the WD cooling for gigayears and has been shown to be very important for explaining several features in WD populations (e.g., Tremblay et al. 2019; Schreiber et al. 2021; Bagnulo & Landstreet 2022; Blatman & Ginzburg 2024).
We therefore improved the WD evolution in our analysis by computing more accurate WD properties via interpolation and extrapolation using the cooling sequences calculated by Bédard et al. (2020), assuming a thin hydrogen atmosphere. We assumed all WDs have thin layers, but checked the impact of the assumed thickness of the hydrogen layer by assuming all have thick layers and obtained very similar magnitudes. The impact was not large enough to affect the results. Provided the WD age and mass, we linearly interpolate–extrapolate through the sequences to obtain the WD properties, such as effective temperature, radius, and luminosity. These sequences were computed assuming solar metallicity, which is higher than assumed in our GC simulations. However, the impact of metallicity on WD evolution regarding the properties in which we are interested is virtually negligible (Renedo et al. 2010). That said, for the purposes of our analysis it is reasonable to use the solar metallicity sequences calculated by Bédard et al. (2020). The main motivation to use such sequences among the few that are available comes from the fact that they were used by Heintz et al. (2022).
@@ 3.2 Initial conditions
We used 197 cluster simulations with a wide range of initial cluster parameters. A full description of the different cluster initial conditions can be found in Hypki et al. (2024). The most important parameters are the following:
  • many different combinations of sizes of clusters, from 550 thousand to 2.6 million initial objects (two multiple stellar populations);
  • both tidally filling and underfilling clusters;
  • both a 10% and 95% initial binary fraction;
  • a Kroupa mass function between 0.08 and 150 M;
  • different Galactocentric distances;
  • different properties of multiple populations.
A big difference with these simulations compared to earlier simulations is that we are using multiple-stellar populations in the framework of the asymptotic giant branch scenario (Calura et al. 2019) in these new simulations, and thus a new path to GC evolution. Instead of having one tidally underfilling dense cluster, we have two populations: a tidally filling (or slightly tidally underfilling) less dense population with a higher maximum initial stellar mass, and a tidally underfilling denser population. This leads to a large number of early escapers from the first population where approximately 30-40% of mass is removed in the first few megayears. These binaries are mostly primordial and undisturbed from dynamical interactions since they escape at an early stage before they have had time to interact with other objects. The second population is denser, but the maximum stellar mass is lower; there is no formation of massive stars in this population. This limit is set to 20 M, which means that there are no supernovae forming BHs from the second population. The previously used initial cluster conditions are quite similar to the second population, while the first population is not as tightly bound and is able to lose a large amount of mass early in the evolution.
The choice of a 10% or 95% binary fraction affects the initial binary properties. For example, in a 10% binary fraction cluster, the largest semi-major axis value is 100 au and the distribution is uniform in log(a). With a 95% binary fraction, we allow binaries to form with larger separations and the initial distribution of semi-major axis and eccentricity is based on Kroupa (1995) and Belloni et al. (2017). Since these binaries are often disrupted by dynamical interactions, a 95% binary fraction has been found to agree well with observations (Belloni et al. 2017, 2018a). For both binary fractions, the initial eccentricity distribution is thermal.